OpenStudy (anonymous):

The slope field to a differential equation is shown below. Which of the following differential equations could be a solution to that differential equation?

OpenStudy (anonymous):

y = 2 + sin x y = 5 - sin x y = 3 + cos x y = 1 - sin x

OpenStudy (anonymous):

I have graphed all of these. I am not sure what to do because none of them match this picture

OpenStudy (anonymous):

Am I supposed to find the derivative of each of these and then graph it?

OpenStudy (anonymous):

I think that it is y=3+cos(x) due to the way that it curves